The Organizing Committee of the 15-th ILAS Conference thanks

Bryan Shader

University of Wyoming, Laramie, WY, USA for organizing the Mini-Symposium on “Combinatorial Matrix Theory".

The Organizing Committee of the 15-th ILAS Conference thanks

Leslie Hogben

Iowa State University, Ames, Iowa, USA for organizing the Mini-Symposium on “Combinatorial Matrix Theory".

The Organizing Committee of the 15-th ILAS Conference thanks

Julio Moro

Universidad Carlos III de Madrid, Leganés, Spain

for organizing the Mini-Symposium on “Eigenproblems: Theory and computation".

The Organizing Committee of the 15-th ILAS Conference thanks

Rosemary A. Renaut

Department of Mathematics and Statistics, Tempe, USA for organizing the Mini-Symposium on

“Implementation and Application issues in regularizing least squares and total least squares".

The Organizing Committee of the 15-th ILAS Conference thanks

María Trigueros

Depto de Matemáticas ITAM, México DF, México for organizing the Mini-Symposium on “Linear Algebra Education".

The Organizing Committee of the 15-th ILAS Conference thanks

Asuman Oktaç

CINVESTAV, México DF, and Pontificia Universidad Católica de

Valparaíso, Chile for organizing the Mini-Symposium on “Linear Algebra Education".

Peter Benner

TU Chemnitz, Fakultät für Mathematik, Chemnitz, Germany

for organizing the Mini-Symposium on “Linear Algebra in Model Reduction".

Chun-Hua Guo

University of Regina, Regina, Canada

for organizing the Mini-Symposium on “Matrix Functions and Matrix Equations".

Peter Lancaster

University of Calgary, Calgary, Canada

for organizing the Mini-Symposium on “Matrix Functions and Matrix Equations".

Peter Butkovic

University of Birmingham, Birmingham, United Kingdom

for organizing the Mini-Symposium on “Max Algebra".

Hans Schneider

University of Wisconsin, Madison, Madison, WI, USA

for organizing the Mini-Symposium on “Max Algebra".

Judith McDonald

Washington State University, Pullman, WA, USA

for organizing the Mini-Symposium on

“Nonnegative and eventually nonnegative matrices".